### v1

# FastGS: Training 3D Gaussian Splatting in 100 Seconds

Abstract
[…]
Figure 1: We propose FastGS, a general acceleration framework for 3D Gaussian Splatting (3DGS) that significantly reduces training time without sacrificing rendering quality. In static scenes, our method completes training on the Tanks & Temples train scene within 100 seconds. Furthermore, our method achieves 2.82× and 2.24× faster training for dynamic and surface reconstruction, respectively.
[…]
1 Introduction
Novel view synthesis (NVS) is a fundamental problem in computer vision and graphics, with broad applications in augmented reality [zhou2018stereo], virtual reality [xu2023vr] and autonomous driving [wu2023mars]. Neural Radiance Field (NeRF) [mildenhall2021nerf] methods model scenes as continuous volumetric functions and render photorealistic views, but require hours of training per scene. Recently, 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] has achieved rendering quality comparable to NeRF while offering significantly faster training and rendering speed. It models 3D scenes via explicit Gaussian primitives and employs a tile-based rasterizer. Benefiting from its efficiency, 3DGS has been successfully applied to a wide range of tasks, including dynamic scene reconstruction, surface reconstruction and simultaneous localization and mapping (SLAM). However, a major bottleneck in its current practical application is the extended training time, which often requires tens of minutes per scene, hindering user-friendly deployment.
A detailed analysis of the vanilla 3DGS training pipeline reveals two primary limitations: (1) its adaptive density control (ADC) of Gaussians often introduces a large number of redundant Gaussians, and (2) inefficiencies in the vanilla 3DGS rendering pipeline. While recent works [mallick2024taming, hanson2025speedy, feng2025flashgs, gui2024balanced] have significantly optimized the rendering pipeline, ADC remains a major area for improvement.
ADC in vanilla 3DGS comprises two main components. The first is Gaussian densification, which clones or splits a Gaussian based on its positional gradient. The second is Gaussian pruning, which removes Gaussians with low opacity or oversized scales. Existing 3DGS acceleration methods [fang2024mini, girish2024eagles, papantonakis2024reducing, lan20253dgs2, wang2024adr, deng2024efficient, mallick2024taming, hanson2025speedy, chen2025dashgaussian] have introduced improvements to ADC. One direction of improvement focuses on designing mechanisms to constrain Gaussian densification, aiming to minimize the growth of redundant Gaussians. For example, Taming-3DGS [mallick2024taming] employs a budget-constrained optimization to control Gaussian growth. Similarly, DashGaussian [chen2025dashgaussian] leverages an adaptive Gaussian primitive budgeting method to maintain continuous densification throughout training. The other direction involves refining the pruning strategy to accelerate training by deleting a greater number of redundant Gaussians. For example, Speedy-Splat [hanson2025speedy] applies a soft pruning strategy during densification and a hard pruning strategy afterward.
One major drawback of these methods is the limited effectiveness of their densification and pruning strategies, which fail to maintain rendering quality while avoiding excessive Gaussian redundancy, resulting in an inefficient representation, as illustrated in Fig. 2. This indicates that their Gaussian control strategies are suboptimal. Based on our observations, some densification and pruning methods [kim2024color, kerbl3Dgaussians, kheradmand20243d, ye2024absgs, chen2025dashgaussian, ali2024trimming, girish2024eagles, salman2024elmgs] do not leverage multi-view consistency, while others [zhang2024pixel, mallick2024taming, hanson2025speedy, hanson2025pup] exploit it suboptimally. Specifically, they enforce multi-view consistency merely through Gaussian-related scores, which we argue is insufficient. It may lead to the excessive growth of redundant Gaussians that provide only marginal improvements to the rendering quality from a few viewpoints while contributing little to others. On one hand, for certain densification methods such as Taming-3DGS [mallick2024taming], Gaussian importance is considered across views. However, it fully relies on Gaussian-related scores rather than their actual contribution to rendering quality, resulting in weak multi-view constraints and leading to redundancy. Moreover, it lacks a dedicated redesign of the pruning strategy. On the other hand, for some pruning methods such as Speedy-Splat [hanson2025speedy], multi-view information is also considered, but it uses the gradients of Gaussian rather than evaluating each Gaussian’s contribution to multi-view rendering quality. This indirect enforcement of multi-view consistency leads to significant degradation in rendering quality.
To address the above issues, we propose FastGS, a new, simple, and general 3DGS acceleration framework, capable of training a scene within 100 seconds while maintaining comparable rendering quality, as shown in Fig. 1. In fact, nearly every Gaussian primitive participates in rendering the same region across multiple viewpoints. Our insight is similar to the concept behind bundle adjustment in traditional 3D reconstruction, where each 3D Gaussian should maintain multi-view consistency. This implies that the 3D Gaussian should enhance rendering quality across multiple views of the same region. Therefore, we introduce a multi-view consistent densification (VCD) strategy, which uses a multi-view reconstruction quality importance score to evaluate whether a Gaussian contributes beneficially to the improvement of multi-view rendering quality. Based on the same idea, we propose a multi-view consistent pruning (VCP) strategy, which removes redundant Gaussians that are useless to multi-view rendering quality. Notably, because VCD and VCP accurately identifies which Gaussians need to be densified or pruned, our method does not require a budget mechanism, making it applicable to other tasks easily. To summarize our contributions,
  • 1.
    New, simple and general framework for 3DGS acceleration that can train a scene within 100 seconds while achieving the comparable rendering quality.
  • 2.
    Efficient densification and pruning strategy strictly controls the addition and removal of each Gaussian based on its contribution to multi-view reconstruction quality, greatly accelerating the training process.
  • 3.
    State-of-the-Art Performance on extensive experiments across Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets. Our method outperforms the state-of-the-art (SOTA) approaches in training efficiency while maintaining comparable rendering quality.
  • 4.
    Strong generality across various tasks, such as dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM. FastGS achieves an average 2–7× training speedup.
Figure 2: Gaussian Count Over Training Iterations. Benefiting from the efficient VCD and VCP strategies, FastGS maintains the smallest number of Gaussians throughout the entire training.
Figure 3: The pipeline of FastGS. (a) We redesign the ADC of the vanilla 3DGS [kerbl3Dgaussians] based on multi-view consistency. To accurately assess the importance of each Gaussian, we sample training views and generate the corresponding per-pixel L1 loss maps. For each sampled view, a multi-view score is computed for each Gaussian by counting the number of high-error pixels within its 2D footprint, which is subsequently used to guide Gaussian densification and pruning. (b) Taming-3DGS [mallick2024taming] primarily computes the importance score based on Gaussian-associated properties across sampled views. (c) Speedy-Splat [hanson2025speedy] computes the Gaussian score by accumulating Gaussian-associated Hessian approximations across all training views. We visualize the densification results from 0.5K to 15K iterations without any pruning on the far left of the figure, and the pruning results after removing Gaussians at the 6K-th iteration on the far right.
[…]
2 Related Work
In this section, we review related works that focus on accelerating both the training and inference of 3DGS [kerbl3Dgaussians].
Gaussian densification.
A key bottleneck of 3DGS is the excessive Gaussians that significantly slow down the training process. Recent works [kim2024color, rota2024revising, mallick2024taming, ye2024absgs, zhang2024pixel, chen2025dashgaussian, kheradmand20243d] attempt to address this issue by refining the densification strategy to control the primitive count. Specifically, 3DGS-MCMC [kheradmand20243d] adopts a Monte Carlo strategy for Gaussian densification. Taming-3DGS [mallick2024taming] employs a budgeting mechanism to control Gaussian growth, primarily computing Gaussian scores based on Gaussian-associated properties. DashGaussian [chen2025dashgaussian] also uses a budgeting method to guide Gaussian densification. Nevertheless, they still require millions of Gaussians to maintain rendering quality, resulting in heavy optimization overhead.
Gaussian Pruning.
Besides modifying densification, some methods [ali2024trimming, salman2024elmgs, hanson2025speedy, hanson2025pup, fan2024lightgaussian, girish2024eagles, fang2024mini, papantonakis2024reducing] achieve acceleration by designing pruning strategies to remove a large number of Gaussians. Some of them design importance scores according to Gaussian properties to guide pruning, discarding less critical Gaussians [ali2024trimming, fan2024lightgaussian, girish2024eagles]. Mini-Splatting [fang2024mini] removes a large number of Gaussians through a simplification strategy based on intersection preserving and sampling. PUP 3DGS [hanson2025pup] and Speedy-Splat [hanson2025speedy] remove Gaussians by computing a Gaussian-associated Hessian approximation across all training views. Nonetheless, some of them fail to fully remove redundant Gaussians, while others remove many Gaussians at the cost of significantly degraded rendering quality.
Other Methods.
Some works focus on optimizing 3DGS rasterization or optimization strategies. Taming-3DGS [mallick2024taming] replaces per-pixel with per-splat parallel backpropagation, which significantly speeds up the optimization process and serves as a strong baseline for subsequent research. StopThePop [radl2024stopthepop], FlashGS [feng2025flashgs], and Speedy-sSplat [hanson2025speedy] use precise tile intersection to reduce Gaussian-tile pairs and accelerate rasterization. 3DGS-LM [hollein20243dgs] replaces Adam with Levenberg-Marquardt for faster convergence, and 3DGS² [lan20253dgs2] achieves near second-order convergence via prioritized per-kernel updates.
3 Background: 3D Gaussian Splatting
3DGS models a scene as an explicit point-based representation composed of a set of anisotropic 3D Gaussians:
$\left\{\mathcal{G}_{i}(\mathbf{x})=\exp\!\left(-\tfrac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\mathbf{\Sigma}_{i_{3D}}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\right)\right\}_{i=1}^{N}.$ (1)
To construct this representation, a sparse point cloud obtained from SfM is used to initialize the positions of Gaussian primitives. Each primitive $\mathcal{G}_{i}$ is parameterized by mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $r_{i}\in\mathbb{R}^{4}$, scale $s_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1)$, and color coefficients $\mathbf{c}_{i}\in\mathbb{R}^{16\times 3}$ represented in view-dependent spherical harmonics (SH). The rotation and scale together define the covariance matrix as
$\mathbf{\Sigma}_{i_{3D}}=R_{i}S_{i}S_{i}^{\top}R_{i}^{\top}.$ (2)
To render from a given camera viewpoint, all 3D Gaussians need to be projected into the 2D image plane. Given a viewing transformation $W$, the covariance matrix in camera coordinates is computed as
$\boldsymbol{\Sigma}^{\prime}_{i_{3D}}=JW\,\boldsymbol{\Sigma}_{i_{3D}}\,J^{\top}W^{\top},$ (3)
where $J$ denotes the Jacobian of the affine approximation of the projective transformation. The projected 3D Gaussian is then approximated as a 2D elliptical Gaussian on the image plane, with covariance $\boldsymbol{\Sigma}_{i_{2D}}$ obtained by marginalizing $\boldsymbol{\Sigma}^{\prime}_{i_{3D}}$ along the viewing direction. Each 2D Gaussian contributes to pixels within its footprint using $\alpha$-blending: given Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$ sorted by depth, the accumulated color of pixel $p$ is computed as
$C(p)=\sum_{i\in N}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}{\left(1-\sigma_{j}\right)}},\quad\alpha_{i}=\sigma_{i}\mathcal{G}_{i}^{\prime}(p),$ (4)
where
$\mathcal{G}^{\prime}_{i}(p)=\exp\Big(-\tfrac{1}{2}(p-\boldsymbol{\mu}_{i_{2D}})^{\top}\boldsymbol{\Sigma}_{i_{2D}}^{-1}(p-\boldsymbol{\mu}_{i_{2D}})\Big),$ (5)
with $\boldsymbol{\mu}_{i_{2D}}$ and $\boldsymbol{\Sigma}_{i_{2D}}$ denoting the mean and covariance of the projected 2D Gaussian, respectively.
4 FastGS
4.1 Overview
The framework of our method is shown in Fig. 3. We initialize 3D Gaussians using SfM point clouds, and train the 3DGS model on multi-view images. The addition and removal of 3D Gaussians are controlled via the proposed multi-view consistent densification and pruning strategies. As illustrated in Fig. 3 and Fig. 3, Taming-3DGS [mallick2024taming] and Speedy-Splat [hanson2025speedy] also estimate Gaussian scores from multiple views, for densification and pruning respectively. However, both rely on Gaussian-associated scores to control the number of Gaussians rather than considering each Gaussian’s contribution to multi-view rendering quality. This suboptimal use of multi-view information results in redundancy for Taming-3DGS [mallick2024taming] and degraded rendering quality for Speedy-Splat [hanson2025speedy] . In contrast, as illustrated in Fig. 3, our method evaluates the importance of each Gaussian based on multi-view reconstruction quality, rather than its properties. Furthermore, our method leverages multi-view consistency for both densification and pruning, which will be detailed in Sections Sec. 4.2 and Sec. 4.3. To further improve rasterization efficiency, Sec. 4.4 introduces the Compact Box (CB) we use.
4.2 Multi-view Consistent Densification
The vanilla 3DGS [kerbl3Dgaussians] densifies Gaussians solely based on the gradient magnitude in the image space, which leads to a large number of redundant Gaussians. Other densification methods [chen2025dashgaussian, ye2024absgs, fang2024mini] also generate millions of Gaussians, leading to inefficiency. We argue that the redundancy arises because these methods fail to rigorously determine from multiple views whether a Gaussian needs densification. As illustrated in Fig. 3, Taming-3DGS [mallick2024taming] considers multi-view consistency during densification. Howerver, it primarily computes the score based on Gaussian-associated properties (e.g., opacity, scale, depth and gradient) rather than their actual contribution to the rendering quality, making it difficult to enforce strict multi-view consistency for a Gaussian. This also leads to redundancy, as visualized on the left of Fig. 3. Moreover, the computation of its score is complex and relatively inefficient. To address these issues, we propose a new, simple densification strategy VCD based on multi-view consistency. As illustrated in Fig. 3, it computes the average number of high-error pixels in each Gaussian’s 2D footprint across sampled views, where high-error pixels are identified solely from the per-pixel L1 loss between the ground truth and the rendering. As shown in the left of Fig. 3, VCD achieves comparable rendering quality with fewer Gaussians, thereby greatly avoiding redundancy. We then detail how VCD is implemented.
Given $K$ camera views $V=\{v^{j}\}_{j=1}^{K}$, randomly sampled from the training views, together with their corresponding ground-truth images $G=\{g^{j}\}_{j=1}^{K}$ and rendered images $R=\{r^{j}\}_{j=1}^{K}$. For each view $v^{j}$, we compute the error between the rendered color $r^{j,c}_{u,v}$ and the ground-truth color $g^{j,c}_{u,v}$ at pixel $(u,v)$:
$e_{u,v}^{j}=\frac{1}{C}\sum_{c=1}^{C}{\left|r_{u,v}^{j,c}-g_{u,v}^{j,c}\right|},$ (6)
where $c\in\{1,2,\dots,C\}$ denotes the color channel. We then construct the loss map $\mathcal{M}^{j}\in\mathbb{R}^{W\times H}$ from the per-pixel errors:
$\mathcal{M}^{j}=\{\,\hat{e}_{u,v}^{j}\,\}_{u=0,v=0}^{W-1,H-1},$ (7)
where each $\hat{e}_{u,v}^{j}$ is obtained via min–max normalization:
$\hat{e}_{u,v}^{j}=\frac{e_{u,v}^{j}-\min\limits_{u,v}(e_{u,v}^{j})}{\max\limits_{u,v}(e_{u,v}^{j})-\min\limits_{u,v}(e_{u,v}^{j})+\varepsilon},$ (8)
with $\varepsilon$ being a small constant for numerical stability. A threshold $\tau$ is then applied to $\mathcal{M}^{j}$ to identify pixel $p_{h}$ with high reconstruction error, forming a mask:
$\mathcal{M}^{j}_{\text{mask}}=\mathbb{I}(\mathcal{M}^{j}>\tau),$ (9)
where pixels $P$ with $\mathcal{M}^{j}_{\text{mask}}(u,v)=1$ indicate regions of poor reconstruction quality.
Next, we need to find the Gaussian primitives associated with these high-error pixels. For each 3D Gaussian primitive $\mathcal{G}_{i}$, we project it onto the 2D image space to obtain its 2D footprint $\Omega_{i}$. We then use an indicator function $\mathbb{I}\big(\mathcal{M}^{j}_{\text{mask}}(p)=1\big)$ to determine whether a pixel has high error. We compute an importance score $s^{+}_{i}$ for each Gaussian primitive, which accumulates the number of high-error pixels contained in the 2D footprint across all sampled views and then averages the accumulated value:
$s_{i}^{+}=\frac{1}{K}\sum_{j=1}^{K}{\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}},$ (10)
where a higher $s^{+}_{i}$ indicates that the Gaussian consistently lies in high-error regions across multiple views, thus suggesting it as a candidate for densification. A Gaussian primitive $\mathcal{G}_{i}$ is selected for densification if its importance score $s^{+}_{i}$ exceeds a predefined threshold $\tau_{+}$, indicating that it contributes significantly to under-reconstructed regions across multiple views. Notably, we can efficiently determine the number of high-error pixels within the 2D footprint directly from the forward pass of the render.
4.3 Multi-view Consistent Pruning
The vanilla 3DGS [kerbl3Dgaussians] removes Gaussians with low opacity or overly large scale but cannot effectively address redundancy. Recent pruning strategies [ali2024trimming, fan2024lightgaussian, fang2024mini] similarly fail to eliminate redundancy and can even significantly degrade rendering quality. In all cases, they do not determine Gaussian redundancy based on multi-view consistency. As illustrated in Fig. 3, Speedy-Splat [hanson2025speedy] computes the pruning score by accumulating Gaussian-associated Hessian approximations across all training views rather than capturing each Gaussian’s true contribution to multi-view rendering quality. Hence, it leads to degraded rendering quality due to its indirect use of multi-view consistency, as visualized on the right of Fig. 3. To remove truly redundant Gaussians, we propose a new, simple pruning strategy VCP based on multi-view consistency. As illustrated in Fig. 3, similar to VCD, it evaluates the score according to each Gaussian’s impact on multi-view reconstruction quality. As shown on the right of Fig. 3, VCP removes a significant number of redundant Gaussians while preserving rendering quality. We then detail how VCP is implemented.
Specifically, for each view $v^{j}\in V$, we compute the photometric loss between the rendered image $r^{j}$ and the corresponding ground-truth image $g^{j}$:
$E_{\mathrm{photo}}^{j}=(1-\lambda)L_{1}^{j}+\lambda(1-L_{\mathrm{SSIM}}^{j}),$ (11)
where $L_{1}^{j}$ and $L_{\mathrm{SSIM}}^{j}$ denote the mean absolute error and the structural similarity loss over the entire image, respectively. Since the photometric loss provides a reliable indicator of reconstruction fidelity, we incorporate it with Eq. 10 to derive the pruning score for each Gaussian primitive $\mathcal{G}_{i}$:
$s_{i}^{-}=\mathcal{N}\left(\sum_{j=1}^{K}{\left(\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}\right)\cdot E_{\mathrm{photo}}^{j}}\right),$ (12)
where $\mathcal{N}(\cdot)$ denotes a min–max normalization function. Here, $s_{i}^{-}$ can be interpreted as a quantitative measure of the contribution of Gaussian primitive $\mathcal{G}_{i}$ to the degradation of the overall rendering quality. A Gaussian primitive $\mathcal{G}_{i}$ is selected for pruning if its importance score $s^{-}_{i}$ exceeds a predefined threshold $\tau_{-}$, indicating that it has relatively low contribution to rendering quality across multiple views.
4.4 Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] uses the 3-sigma rule to obtain 2D ellipses, generating many Gaussian–tile pairs that introduce computational redundancy and reduce rendering efficiency. Speedy-Splat [hanson2025speedy] partially addresses this with precise tile intersect, yet we observe that some 2D Gaussians still have negligible impact on pixels in certain tiles. To further reduce unnecessary pairs, we propose Compact Box, which prunes Gaussian–tile pairs with minimal contribution based on their Mahalanobis distance from the Gaussian center, thereby accelerating rendering while maintaining quality. Details are provided in the supplementary material.
4.5 Optimization
Same as the vanilla 3DGS [kerbl3Dgaussians], we optimize the learnable parameters with respect to the L1 loss over rendered pixel colors, combined with the SSIM term [SSIM] $\mathcal{L}_{\text{SSIM}}$. The total supervision is defined as:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda(1-\mathcal{L}_{\text{SSIM}}).$ (13)
Table 1: Quantitative comparisons with existing 3DGS fast optimization methods. With FastGS, the training of 3DGS can be completed within 100 seconds on average, while achieving comparable rendering quality to the other methods. Best results are marked as best score, second best score and third best score. Time is reported in minutes.
Method Mip-NeRF 360 [barron2022mip] Deep Blending [hedman2018deep] Tanks & Temples [knapitsch2017tanks]
[…]
3DGS [kerbl3Dgaussians] 20.93 27.53 0.813 0.221 2.63M 146 19.77 29.71 0.903 0.241 2.46M 158 11.34 23.71 0.850 0.170 1.57M 195
3DGS-accel [kerbl3Dgaussians, mallick2024taming] 10.94 27.52 0.813 0.221 2.63M 188 8.87 29.74 0.903 0.242 2.46M 230 6.96 23.85 0.850 0.170 1.57M 252
Mini-Splatting [fang2024mini] 17.71 27.33 0.821 0.217 0.53M 564 13.35 29.99 0.907 0.244 0.56M 624 9.06 23.46 0.844 0.181 0.30M 756
Speedy-splat [hanson2025speedy] 13.38 26.91 0.781 0.295 0.30M 552 10.75 29.42 0.898 0.272 0.25M 664 6.32 23.38 0.816 0.242 0.18M 691
Taming-3DGS [mallick2024taming] 5.36 27.48 0.794 0.261 0.68M 221 3.06 29.50 0.894 0.278 0.29M 352 2.71 23.89 0.833 0.214 0.32M 379
DashGaussian [chen2025dashgaussian] 6.38 27.73 0.817 0.217 2.40M 154 4.16 29.65 0.906 0.246 1.94M 208 4.28 24.00 0.853 0.178 1.21M 240
FastGS-20k (Ours) 1.32 27.35 0.793 0.268 0.39M 568 0.88 29.93 0.901 0.274 0.22M 687 0.95 23.81 0.835 0.216 0.25M 632
FastGS (Ours) 1.92 27.54 0.796 0.264 0.38M 584 1.28 30.03 0.901 0.270 0.22M 714 1.32 24.15 0.839 0.210 0.24M 655
FastGS-Big (Ours) 3.59 27.93 0.820 0.216 1.16M 472 2.00 30.12 0.907 0.242 0.65M 607 2.03 24.39 0.855 0.175 0.54M 569
Figure 4: Qualitative results of Tab. 1. We present qualitative results on the “kitchen” and “stump” scenes from the Mip-NeRF 360 dataset [barron2022mip], the “playroom” scene from the Deep Blending dataset [deng2022depth], and the “truck” scene from the Tanks&Temples dataset [knapitsch2017tanks].
5 Experiments
5.1 Experimental Setup
Datasets. Same as vanilla 3DGS [kerbl3Dgaussians], we conduct experiments on three real-world datasets: Mip-NeRF 360 [barron2022mip], Deep-Blending [hedman2018deep], and Tanks & Temples [knapitsch2017tanks]. For other challenging tasks, we evaluate dynamic scene reconstruction on D-NeRF [pumarola2021d], NeRF-DS [yan2023nerf], and Neu3D [li2022neural] datasets. Tanks & Temples [knapitsch2017tanks], LLFF [mildenhall2019local], BungeeNeRF [xiangli2022bungeenerf] and Replica [replica19arxiv] are respectively used for surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM.
[…]
Method Mip-NeRF 360 [barron2022mip] Deep Blending [hedman2018deep] Tanks & Temples [knapitsch2017tanks]
[…]
3DGS [kerbl3Dgaussians] 20.93 27.53 0.813 0.221 2.63M 146 19.77 29.71 0.903 0.241 2.46M 158 11.34 23.71 0.850 0.170 1.57M 195
+Ours 9.18 27.54 0.796 0.264 0.39M 504 9.60 30.09 0.901 0.270 0.22M 618 4.93 24.24 0.840 0.209 0.24M 586
Mip-Splatting [yu2024mip] 26.20 27.89 0.837 0.176 3.98M 224 23.94 29.35 0.899 0.241 3.48M 219 14.57 23.77 0.856 0.158 2.36M 300
+Ours 2.98 27.95 0.828 0.208 0.83M 606 1.74 29.68 0.899 0.274 0.20M 698 2.01 24.18 0.843 0.200 0.36M 729
Scaffold-GS [lu2024scaffold] 18.37 27.70 0.812 0.226 0.57M 194 13.31 30.09 0.905 0.256 0.18M 307 9.83 24.09 0.851 0.175 0.26M 261
+Ours 5.06 27.68 0.809 0.220 0.30M 281 2.82 30.00 0.900 0.267 0.08M 423 3.77 24.15 0.849 0.180 0.14M 332
3DGS-accel [kerbl3Dgaussians, mallick2024taming] 10.94 27.52 0.813 0.221 2.63M 188 8.87 29.74 0.903 0.242 2.46M 230 6.96 23.85 0.850 0.170 1.57M 252
+Ours 1.92 27.54 0.796 0.264 0.38M 584 1.28 30.03 0.901 0.270 0.22M 714 1.32 24.15 0.839 0.210 0.24M 655
[…]
NeRF-DS [yan2023nerf] Deformable-3DGS [yang2024deformable] 7.86 23.83 0.851 0.180 0.13M 103
+Ours 2.75 23.90 0.854 0.176 0.03M 484
Neu3D [li2022neural] Deformable-3DGS [yang2024deformable] 25.11 27.78 0.902 0.158 0.20M 73
+Ours 8.70 30.25 0.919 0.140 0.08M 172
Metrics. To evaluate the performance, we report commonly used metrics for novel view rendering quality, including PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable]. In addition, training efficiency and model compactness are assessed by reporting the total training time (in minutes), the final number of Gaussians, and the rendering speed (FPS).
Implementation Details. All methods, including ours and the other compared approaches, are trained for 30K iterations using the Adam [kingma2014adam] optimizer. For our approach, we set $K=10$ and $\lambda=0.2$ in all experiments. In the base setting, densification is performed every 500 iterations until the 15,000-th iteration. To ensure fairness, all experiments are conducted with an NVIDIA RTX 4090 GPU, and all comparison methods are implemented using their official code. The default configuration of FastGS is 3DGS-accel [kerbl3Dgaussians, mallick2024taming] with the proposed VCD, VCP, and CB. To achieve extreme training acceleration, the rendering quality of our method is not the highest. Therefore, we provide a variant, FastGS_\_Big, which achieves both the highest rendering quality and the fastest training speed, where densification is executed once every 100 iterations. Further details are in the supplementary material.
Figure 5: Visualization results on four representative tasks. We present the rendered results of the “coffee-martini”, “Ignatius”, “fern”, and “bilbao” scenes from the Neu3D [li2022neural], Tanks and Temples [knapitsch2017tanks], LLFF [mildenhall2019local], and BungeeNeRF [xiangli2022bungeenerf] datasets, respectively.
5.2 Comparison with Fast Optimization Methods
Baselines. We adopt 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as our baseline which preserves the original 3DGS [kerbl3Dgaussians] pipeline and incorporates the efficient per-splat parallel backpropagation proposed by Taming-3DGS [mallick2024taming]. Beyond this baseline, we further conduct comparisons against several recent SOTA acceleration techniques, including DashGaussian [chen2025dashgaussian], Mini-Splatting [fang2024mini], Speedy-Splat [hanson2025speedy], and Taming-3DGS [mallick2024taming], which represent complementary advances in accelerating training from different perspectives.
Quantitative Results. As shown in Tab. 1, FastGS achieves the fastest training with comparable rendering quality. On average, a scene can be trained within 100 seconds, and the fastest case takes only 77 seconds. This benefit comes from our effective multi-view consistent densification and pruning strategies. Speedy-Splat [hanson2025speedy] fails to fully consider multi-view consistency during pruning, leading to a significant drop in rendering quality. Taming3DGS [mallick2024taming] applies weak multi-view consistency constraints, resulting in excessive Gaussian growth and slower training. The current SOTA, DashGaussian [chen2025dashgaussian], achieves high rendering quality yet lacks an effective pruning strategy to remove redundant Gaussians. Its scene optimization still retains several million Gaussians, limiting training speed. In contrast, our variant FastGS-Big surpasses DashGaussian [chen2025dashgaussian] by more than 0.2 dB in rendering quality, reduces training time by over 43.7%, and cuts the number of Gaussians by more than half. These results demonstrate the superiority of our multi-view consistent densification and pruning strategies.
We visualize the results of Tab. 1 in Fig. 4, providing a detailed comparison of reconstructed scenes across different methods. Our method FastGS achieves comparable rendering quality without losing any details. In addition, our variant FastGS-Big reconstructs scenes more accurately, with more precise geometric structures under the multi-view consistency constraint.
[…]
Tanks & Temples [knapitsch2017tanks] PGSR [chen2024pgsr] 32.92 24.20 0.857 0.149 1.56M 87 0.57
+Ours 15.73 24.37 0.845 0.190 0.40M 210 0.55
Mip-NeRF 360 [barron2022mip] PGSR [chen2024pgsr] 74.70 27.22 0.832 0.183 3.79M 45 -
+Ours 11.68 27.23 0.813 0.240 0.49M 202 -
5.3 Generality of FastGS
Baselines. Deformable-3DGS [yang2024deformable], PGSR [chen2024pgsr], DropGaussian [park2025dropgaussian], OctreeGS [ren2024octree] and Photo-SLAM [hhuang2024photoslam] are selected as the backbones for dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM, respectively.
[…]
Enhancing Various Tasks. We further test several SOTA methods combined with our framework across these tasks. As shown in Tab. 3 and Tab. 4, our method improves the training speed of all baselines by 2-7× while preserving rendering quality. We visualize some results in Fig. 5. This improvement demonstrates the strong generality of our approach. We argue that this benefits from the multi-view consistency, which is fundamental to various reconstruction tasks. More results are provided in the supplementary material and we will open-source our code.
5.4 Ablation Study
We adopt 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as our baseline and systematically evaluate the contribution of each proposed module. By adding each component individually, we analyze its impact on both reconstruction quality and training efficiency, providing insights into the effectiveness of our design.
Multi-View Consistent Densification. We first evaluate the effect of the densification strategy VCD. As shown in  Tab. 5, VCD achieves nearly 3× faster training without any loss in reconstruction quality. This is because our Gaussian addition is guided by stricter multi-view consistency, which prevents the addition of redundant Gaussians.  Tab. 5 further validates this by showing that with VCD, the number of Gaussians is reduced by 80%. This ablation study demonstrates the effectiveness of VCD for accelerating training.
Table 5: Ablation studies over the proposed methods of FastGS. Experiments are performed on the Mip-NeRF 360 dataset [barron2022mip] with 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as the baseline.
[…]
3DGS-accel 10.94 27.52 0.813 0.221 2.63M 188
+VCD. 3.60 27.58 0.798 0.260 0.51M 228
+VCP. 5.60 27.53 0.812 0.228 1.87M 292
+CB. 10.08 27.51 0.810 0.219 2.76M 325
Full 1.92 27.54 0.796 0.264 0.38M 584
Multi-View Consistent Pruning. Next, we evaluate the effectiveness of VCP. As shown in  Tab. 5, adding VCP shortens the training time by 48.8% and reduces the number of Gaussians by 28.9%, without sacrificing rendering quality. This is because our method effectively removes redundant Gaussians while preserving those critical for scene reconstruction. The strict multi-view consistency evaluation for each deleted Gaussian ensures this effectiveness, demonstrating that VCP is highly effective.
Compact Box. Finally, we evaluate the effectiveness of Compact Box. As shown in  Tab. 5, adding CB shortens the training time by 7.8% while achieving the comparable rendering quality. This demonstrates that CB can accelerate training without degrading reconstruction quality.
5.5 Discussions and Limitations
We identify two additional directions to further accelerate training. (1) Our current bottleneck lies in the need for 30k optimization steps. Designing faster-converging optimization strategies that require fewer steps could significantly improve training speed. (2) Another bottleneck is that training starts from scratch with COLMAP initialization. Developing a general feed-forward NVS model and fine-tuning on its results could also greatly shorten training time.
6 Conclusion
This paper introduces a novel, simple, and general acceleration framework FastGS. We propose multi-view consistent Gaussian densification and pruning strategies that prevent redundant Gaussians and significantly accelerate 3DGS training. Extensive experiments demonstrate the effectiveness of our view-consistency design. Our method achieves the fastest training speed among all SOTA methods while maintaining comparable rendering quality. The results also demonstrate strong generality of FastGS, greatly reducing training time across various tasks such as dynamic scene reconstruction, surface reconstruction, and so on.